OpenStudy (anonymous):

Find the probability of getting 2 tails when 3 fair coins are tossed

OpenStudy (raden):

all events = 2^3 = 8

OpenStudy (raden):

the events 2 tails = 3!/2!1! = 3

OpenStudy (zehanz):

So it has to be: TTH, or THT, or HTT. The probabilities of H and T are both 1/2 (fair coins) Next step?

OpenStudy (raden):

so, the probility = ....

OpenStudy (raden):

*probability

OpenStudy (zehanz):

Proability of TTH is (1/2) * (1/2) * (1/2) = Same for the other outcomes, so ...

OpenStudy (zehanz):

Hey RadEn, I like the different methods we use to solve this ;)